OpenStudy (anonymous):

The integral of x/(sqrt(3-x^4)dx

OpenStudy (anonymous):

\[\int\limits_{}^{} x/\sqrt{3-x^4}dx\]

OpenStudy (anonymous):

Well, let u=x^2 Then du=2xdx (1/2)du=xdx Your integral becomes: \[\frac{1}{2}\int\limits \frac{u}{\sqrt{3-u^2}}du\] From here you can use trig sub. Do you need me to take it further?

OpenStudy (anonymous):

I thought I wouldn't be able to use trig sub because of the 3...

OpenStudy (anonymous):

Of course you can. It doesn't have to be a 1. Or a perfect square for that matter. You have: \[\sqrt{(\sqrt{3})^2-u^2}\] Use sin(theta) for your trig sub.

OpenStudy (anonymous):

Oh, okay I thought you had to have a number that was a perfect square :)

OpenStudy (anonymous):

Oh and quick question, how does the x on top become u?

OpenStudy (anonymous):

It shouldn't be. I just mistyped it.

OpenStudy (anonymous):

It should be just du.

OpenStudy (anonymous):

Tell me if you get stuck :P

OpenStudy (anonymous):

So the x on top gets cancelled out?

OpenStudy (anonymous):

Yeah, when you replace it with du. :)

OpenStudy (anonymous):

Oh yea, okay I see once I wrote it down :) Thank you!